# Simplifying Traffic Anomaly Detection with Video Foundation Models

### v1

@@ Abstract
Recent methods for ego-centric Traffic Anomaly Detection (TAD) often rely on complex multi-stage or multi-representation fusion architectures, yet it remains unclear whether such complexity is necessary. Recent findings in visual perception suggest that foundation models, enabled by advanced pre-training, allow simple yet flexible architectures to outperform specialized designs. Therefore, in this work, we investigate an architecturally simple encoder-only approach using plain Video Vision Transformers (Video ViTs) and study how pre-training enables strong TAD performance. We find that: (i) strong pre-training enables simple encoder-only models to match or even surpass the performance of specialized state-of-the-art TAD methods, while also being significantly more efficient; (ii) although weakly- and fully-supervised pre-training are advantageous on standard benchmarks, we find them less effective for TAD. Instead, self-supervised Masked Video Modeling (MVM) provides the strongest signal; and (iii) Domain-Adaptive Pre-Training (DAPT) on unlabeled driving videos further improves downstream performance, without requiring anomalous examples. Our findings highlight the importance of pre-training and show that effective, efficient, and scalable TAD models can be built with minimal architectural complexity. We release our code, domain-adapted encoders, and fine-tuned models to support future work: https://github.com/tue-mps/simple-tad.
@@ 1 Introduction
Figure 1: Traffic Anomaly Detection (TAD) performance on DADA-2000 [13]. Simple encoder-only models with strong pre-training (blue) are faster and more accurate than recent multi-component architectures (orange and red), see Tabs. 2 and 3.
@@
Traffic risk estimation is fundamental to safe driving, as failures to anticipate danger can lead to life-threatening consequences. Autonomous vehicles must therefore assess potential hazards in real time, even under uncertain, dynamic, and unfamiliar conditions. A common formulation for this problem is the ego-centric traffic anomaly detection (TAD) task [48, 13], which aims to identify abnormal or dangerous events in a video stream captured by a vehicle‐mounted camera. Analyzing the top-performing TAD methods [51, 24, 32, 25], we find that they rely on specialized, complex architectures illustrated in Fig. 2(b) and (c): two-stage approaches [51, 34, 10], which combine a vision encoder with a temporal component, and multi-representation fusion approaches [24, 32, 25], which fuse additional representations, often generated by separate deep neural networks or model-based algorithms. While these complex designs have improved performance, their impact on efficiency has not been evaluated, even though this is crucial for TAD, where rapid detection is needed to enable timely action and prevent accidents.
@@
Figure 2: Types of model architectures for TAD: simple encoder-only model (a), two-stage design (b), and multi-representation fusion architecture (c).
@@
Considering that TAD is, in essence, a binary classification task on video, we turn to recent methods for general video classification. On standard video classification benchmarks [19, 15, 8, 36], Video Foundation Models (ViFMs) achieve state-of-the-art performance, predominantly Video Vision Transformers (ViTs) [11, 1], which rely on large-scale self- and weakly-supervised pre-training to learn expressive and transferable spatiotemporal representations, rather than on architectural inductive biases. Indeed, recent work in related visual perception tasks shows that strong pre-training reduces the need for downstream task-specific components [21, 41, 20]. We hypothesize that the same applies to TAD, such that a simple ViT-based ViFM can be effectively applied to this task and match or even outperform complex architectures. Since TAD depends on motion understanding, we investigate whether pre-training strategies that capture spatio-temporal structure are particularly effective.
@@
To test our hypotheses, we evaluate multiple ViT-based ViFMs, sharing the same plain video ViT architecture but different pre-training, on two TAD datasets, DoTA [48] and DADA-2000 [13]. We adopt an encoder-only design, with a single linear layer on top of the Video ViT model, as illustrated in Fig. 2(a). While prior work has attempted the encoder-only design [10], it was found to be inferior [24, 32, 34]. We revisit this design using stronger pre-training. Similar to prior work, we assess in-domain and out-of-domain generalization, and unlike prior work, we also assess computational efficiency of the models, and compare to the best-performing specialized TAD methods.
@@
We confirm our hypothesis by showing that strong pre-training enables a plain video ViT, used as an encoder-only model for TAD, to match and even surpass state-of-the-art methods while also being significantly more efficient, as shown in Fig. 1 and Tab. 2. Interestingly, performance on standard video classification benchmarks does not correlate with TAD. Our comparison of pre-trained models shows that weak supervision from language and full supervision from class labels are effective on standard benchmarks but less so for TAD, likely because they promote appearance-focused features that generalize poorly to anomalous motion [45]. In contrast, self-supervised learning with Masked Video Modeling (MVM), which trains the model to reconstruct missing spatio-temporal regions using both spatial structure and temporal continuity, proves most effective for TAD. Models pre-trained with this objective achieve state-of-the-art performance at their respective model size, as shown in Tab. 3.
@@
Next, motivated by the scarcity of labeled data for TAD and the abundance of unlabeled driving videos, we explore whether we can leverage self-supervised learning to better adapt an off-the-shelf ViFM to the downstream domain. Specifically, we apply Domain-Adaptive Pre-Training (DAPT) [16] using the Video Masked Autoencoding (MAE) [38] approach. We find that MAE-based DAPT, even when applied at relatively small scale compared to the preceding generic pre-training, significantly boosts the performance even further, particularly for smaller models, as shown in Fig. 3. Importantly, we find that including abnormal driving examples for DAPT is not necessary, as shown in Tab. 4, which is valuable given the difficulty of collecting them at scale.
@@
    We show that a plain encoder-only Video ViT, when equipped with strong pre-training, outperforms all prior specialized architectures for TAD, while also being significantly more efficient.
@@ 2.1 Traffic Anomaly Detection
Traffic Anomaly Detection (TAD) is typically framed as a binary video classification task, where the goal is to detect potentially dangerous or abnormal events in traffic scenarios from the egocentric viewpoint of a vehicle-mounted camera. While related to the broader field of Video Anomaly Detection (VAD), which commonly targets static-camera surveillance settings, TAD poses unique challenges due to ego-motion, frequent occlusions, and the dynamic interaction of agents in complex driving environments.
Before the availability of annotated datasets for TAD, earlier approaches often relied on unsupervised reconstruction or prediction of video frames to flag anomalies using temporal autoencoders [29, 43], future frame prediction with spatial, temporal, and adversarial losses [26], or spatiotemporal tubelet modeling with ensemble scoring [48]. Some works also explored the use of synthetic data for training [35, 22].
The introduction of large-scale driving anomaly datasets [48, 13] with comprehensive annotations has enabled more active development of fully-supervised methods and led to substantial improvements in detection performance. As shown in Fig. 2, we categorize TAD methods into three classes based on their architectural complexity, which we detail below.
@@
Encoder-only design, (Fig. 2, a).  Since TAD is a binary classification task, a minimal solution consists of a feature encoder followed by a linear classifier, without additional task-specific modules. Prior work [10] shows that such designs can be effective in related tasks, where R(2+1)D [39] and ViViT [1] demonstrate considerably strong performance. However, several recent studies report underwhelming results for encoder-only ablation variants of their methods [24, 32]. Our evaluation of R(2+1)D and ViViT models on standard TAD benchmarks (see Tab. 2) reveals a substantial performance gap between these encoder-only models and current state-of-the-art methods.
@@
Two-stage design, (Fig. 2, b).  Two-stage methods combine a visual encoder with a separate temporal module. VidNeXt [10] pairs a ConvNeXt [27] backbone with a non-stationary transformer (NST) to model both stable and dynamic temporal patterns, evaluating and introducing a new dataset CycleCrash [10] for the related task of collision prediction. Its ablations, ConvNeXt+VT and ResNet+NST, also yield strong results. MOVAD [34] uses a VideoSwin Transformer [28] as a short-term memory encoder over several frames, followed by an LSTM-based long-term module, achieving state-of-the-art performance for TAD.
@@
Multi-representation fusion design, (Fig. 2, c).  Fusion-based models, which currently report state-of-the-art performance in TAD, explicitly combine multiple information sources. TTHF [24] augments the CLIP [33] framework with a high-frequency motion encoder and a cross-modal fusion module to align motion features with textual prompts. PromptTAD [32] extends MOVAD [34] by incorporating bounding box prompts via instance- and relation-level attention, enhancing object-centric anomaly localization. ISCRTAD [25] integrates agent features (e.g., appearance, trajectory, depth) using graph-based modeling and fuses them with scene context through contrastive multimodal alignment for robust anomaly detection.
@@
In this work, we follow the simple encoder-only design and explore whether strong pre-training can compensate for the lack of task-specific architectural inductive biases. We demonstrate that, when equipped with rich priors from large-scale self-supervised pre-training, such models can achieve state-of-the-art performance, while remaining architecturally simple and highly efficient.
@@ 2.2 Video Foundation Models
2021 ViViT [1] Stage 1 FSL Classification Class labels
2022 VideoMAE [38] Stage 1 SSL Masked Autoencoder Video frame pixels
2023 MVD [45] Stage 1 SSL Masked feature distillation High-level features of VideoMAE and ImageMAE teachers
2023 VideoMAE2 [44] Stage 1 SSL Dual MAE Video frame pixels
Stage 2 FSL Classification Class labels
Stage 3 FSL Logit distillation Logits of larger VideoMAE2 after stage 2
2024 InternVideo2 [46] Stage 1 WSL+SSL Unmasked feature distillation Features of VideoMAE2 and a vision-language encoder
Stage 2 WSL Feature distillation + Contrastive Features of audio and text encoders + video, text, audio
Stage 3 WSL+SSL Feature distillation Features of InternVideo2 after stage 2 across multiple depths
@@
While early 3D CNN-based architectures can be referred to as foundation models [30], the term foundation model today commonly refers to Transformer-based [40] models that leverage large-scale pre-training. For vision, these models typically adopt the Vision Transformer [11] (ViT) architecture.
Unlike the convolutional architecture, which embeds strong spatial and temporal inductive biases, ViTs rely on learning such priors directly from data during pre-training [11]. The quality and scale of this pre-training directly influence their effectiveness in downstream tasks [50]. Pre-training methods vary in supervision type and scalability. Fully-supervised approaches rely on manually annotated labels, providing precise semantic guidance but limited scalability. Weakly-supervised methods, such as CLIP [33], use natural language or metadata as training signals. Though less curated, they offer rich semantic structure and broader concept coverage from web-scale data. Self-supervised learning (SSL) methods, including masked modeling [17, 38], learn from the data itself without any annotations, enabling large-scale pre-training and highly transferable representations. Unlike weak supervision, which relies on sparse and often noisy text labels [33], SSL provides denser and more unbiased training signals [6, 17].
First Video ViTs exploited supervised classification as their pre-training method. ViViT [1] adopted the ViT architecture to video by introducing spatiotemporal 3D cubes called tubelets instead of 2D patches used for images. While this demonstrated that attention-based models can handle video inputs, the method struggled to balance accuracy and efficiency. Subsequent models like TimeSformer [4], MViT [12], and VideoSwin [28] focused on improving efficiency.
VideoMAE [38] adopted masked autoencoding (MAE), a type of Masked Video Modeling (MVM), as an effective and efficient self-supervised pre-training strategy for plain video ViTs. Its tube masking, when applied to a large fraction of input patches, forces the model to infer spatiotemporal structure from limited visible content. This approach yielded strong results while maintaining architectural simplicity and has since inspired a series of ViT-based Video Foundation Models (ViFMs) that employ self-supervised pre-training [45, 44, 46].
@@
While other architectures such as recurrent, hybrid, and state-space models are active research areas [23, 47, 31], at the moment, ViT-based ViFMs are arguably the dominant paradigm due to their strong performance, scalability, versatile pre-training strategies, and widespread optimization support (e.g., FlashAttention [9], optimized libraries, hardware acceleration) designed around the plain Transformer [11] architecture. These qualities, along with the growing availability of pre-trained Video ViTs, make them particularly promising for tasks like traffic anomaly detection (TAD), where generalization, robustness, and efficiency are critical.
@@
To our knowledge, we are the first to research which type of video pre-training is most effective for the TAD task, and hypothesize that MVM, with its emphasis on learning patch-dense and temporally-aware representations, is particularly well-suited for this task.
@@ 3 Methodology
We fine-tune multiple ViT-based Video Foundation Models (ViFMs) for TAD and evaluate their performance against recent specialized TAD methods. We follow the encoder-only design and attach a single linear classification head to the output of the final encoder layer. This minimal design ensures that performance primarily reflects the effectiveness of the ViFM backbone in capturing patterns relevant for traffic anomaly detection. We investigate (i) whether a simple Video ViT model, pre-trained at scale, can achieve state-of-the-art performance on TAD, (ii) whether better general ViFMs are also better for TAD, and what type of pre-training is more effective (iii) and, finally, whether small-scale domain-adaptive pre-training (DAPT) is feasible and effective for adapting Video ViTs to the driving domain.

### v2

@@ Abstract
Recent methods for ego-centric Traffic Anomaly Detection (TAD) often rely on complex multi-stage or multi-representation fusion architectures, yet it remains unclear whether such complexity is necessary. Recent findings in visual perception suggest that foundation models, enabled by advanced pre-training, allow simple yet flexible architectures to outperform specialized designs. Therefore, in this work, we investigate an architecturally simple encoder-only approach using plain Video Vision Transformers (Video ViTs) and study how pre-training enables strong TAD performance. We find that: (i) advanced pre-training enables simple encoder-only models to match or even surpass the performance of specialized state-of-the-art TAD methods, while also being significantly more efficient; (ii) although weakly- and fully-supervised pre-training are advantageous on standard benchmarks, we find them less effective for TAD. Instead, self-supervised Masked Video Modeling (MVM) provides the strongest signal; and (iii) Domain-Adaptive Pre-Training (DAPT) on unlabeled driving videos further improves downstream performance, without requiring anomalous examples. Our findings highlight the importance of pre-training and show that effective, efficient, and scalable TAD models can be built with minimal architectural complexity. We release our code, domain-adapted encoders, and fine-tuned models to support future work: https://github.com/tue-mps/simple-tad.
@@ 1 Introduction
Figure 1: Traffic Anomaly Detection (TAD) performance on DADA-2000 [10]. Simple encoder-only models with advanced pre-training (blue) are faster and more accurate than recent multi-component architectures (orange and red), see Tabs. 2 and 3.
@@
Traffic risk estimation is fundamental to safe driving, as failures to anticipate danger can lead to life-threatening consequences. Autonomous vehicles must therefore assess potential hazards in real time, even under uncertain, dynamic, and unfamiliar conditions. A common formulation for this problem is the ego-centric traffic anomaly detection (TAD) task [54, 10], which aims to identify abnormal or dangerous events in a video stream captured by a vehicle‐mounted camera. Analyzing the top-performing TAD methods [57, 37, 8, 24, 35, 25], we find that they rely on specialized, complex architectures illustrated in Fig. 2, b and c: two-stage approaches [57, 37, 8], which combine a vision encoder with a temporal component, and multi-representation fusion approaches [24, 35, 25], which fuse additional representations, often generated by separate deep neural networks or model-based algorithms. While these complex designs have improved performance, their impact on efficiency has not been evaluated, even though this is crucial for TAD, where rapid detection is needed to enable timely action and prevent accidents.
@@
Figure 2: Types of model architectures for TAD: simple encoder-only (a), two-stage (b), and multi-representation fusion (c).
@@
Considering that TAD is, in essence, a binary classification task on video, we turn to recent methods for general video classification. On standard video classification benchmarks [17, 12, 6, 40], Video Foundation Models (ViFMs) achieve state-of-the-art performance, predominantly Video Vision Transformers (ViTs) [9, 1], which rely on large-scale self- and weakly-supervised pre-training to learn expressive and transferable spatio-temporal representations, rather than on architectural inductive biases. Recent work in related visual perception tasks shows that advanced pre-training reduces the need for downstream task-specific components [19, 47, 18]. We hypothesize that the same applies to TAD, such that a simple ViT-based ViFM can be applied to this task and match or even outperform complex architectures. Since TAD depends on motion understanding, we investigate whether pre-training strategies that capture spatio-temporal structure are particularly effective.
@@
To test our hypotheses, we evaluate multiple ViT-based ViFMs, sharing the same plain Video ViT architecture but different pre-training, on two TAD datasets: DoTA [54] and DADA-2000 [10]. We adopt an encoder-only design, with a single linear layer on top of the Video ViT model, as illustrated in Fig. 2, a. While prior work has attempted the encoder-only design [8], it was found to be inferior [24, 35, 37]. We revisit this design using stronger pre-training. Similar to prior work, we assess in-domain and out-of-domain generalization, and unlike prior work, we assess the computational efficiency of the models and compare to the best-performing specialized TAD methods.
@@
We confirm our hypothesis by showing that advanced pre-training enables a plain Video ViT, used as the encoder in an encoder-only design for TAD, to match and even surpass state-of-the-art methods while also being significantly more efficient, as shown in Fig. 1 and Tab. 2. Interestingly, performance on standard video classification benchmarks does not correlate with TAD. Our comparison of pre-trained models shows that weak supervision from language and full supervision from class labels are effective on standard benchmarks but less so for TAD, likely because they promote appearance-centric features and semantics that generalize poorly to anomalous motion [51]. In contrast, self-supervised learning with Masked Video Modeling (MVM), which trains the model to reconstruct missing spatio-temporal regions using both spatial structure and temporal continuity, proves most effective for TAD. Models pre-trained with this objective achieve state-of-the-art performance at their respective model size, as shown in Tab. 3.
@@
Next, motivated by the scarcity of labeled data for TAD and the abundance of unlabeled driving videos, we explore whether we can leverage self-supervised learning to better adapt an off-the-shelf ViFM to the downstream driving domain. Specifically, we apply Domain-Adaptive Pre-Training (DAPT) [13] using the Video Masked Autoencoding (MAE) [44] approach on unlabeled driving videos. We find that MAE-based DAPT, even when applied at a relatively small scale compared to the preceding generic pre-training, significantly boosts the performance even further, particularly for smaller models, as shown in Fig. 3. Importantly, we find that including abnormal driving examples for DAPT is not necessary, as shown in Tab. 4, which is valuable given the difficulty of collecting them at scale.
@@
    We show that a plain encoder-only Video ViT, when equipped with advanced pre-training, outperforms all prior specialized architectures for TAD, while also being significantly more efficient.
@@ 2.1 Traffic Anomaly Detection
Traffic Anomaly Detection (TAD) is typically framed as a binary video classification task, where the goal is to detect potentially dangerous or abnormal events in traffic scenarios from the egocentric viewpoint of a vehicle-mounted camera. While related to the broader field of Video Anomaly Detection, which commonly targets static-camera surveillance settings, TAD poses unique challenges due to ego-motion, frequent occlusions, and dynamic interactions of agents in complex driving environments. Before the availability of labeled datasets for TAD, earlier approaches often relied on unsupervised reconstruction or prediction of video frames to flag anomalies using temporal autoencoders [30, 49], future frame prediction with spatial, temporal, and adversarial losses [26], or spatio-temporal tubelet modeling with ensemble scoring [54]. Some works also explored the use of synthetic data [39, 20]. The introduction of large-scale driving anomaly datasets [54, 10] with comprehensive annotations has enabled more active development of fully-supervised methods and led to substantial improvements in detection performance. As shown in Fig. 2, we categorize TAD methods into three classes based on their architectural complexity, which we detail below:
@@
Encoder-only design (Fig. 2, a).  Since TAD is a binary classification task, a minimal solution consists of a feature encoder followed by a linear classifier, without any task-specific modules. Prior work [8] shows that such designs can be effective in related tasks, where R(2+1)D [45] and ViViT [1] demonstrate considerably strong performance. However, recent studies report underwhelming results for encoder-only ablation variants of their methods [24, 35, 37].
@@
Two-stage design (Fig. 2, b).  Two-stage methods combine a visual encoder with a separate temporal module. VidNeXt [8] pairs a ConvNeXt [27] backbone with a non-stationary transformer (NST) to model both stable and dynamic temporal patterns, evaluating and introducing a new dataset CycleCrash [8] for the related task of collision prediction. Its ablations, ConvNeXt+VT and ResNet+NST, also yield strong results. MOVAD [37] uses a VideoSwin Transformer [28] as a short-term memory encoder over several frames, followed by an LSTM-based long-term module, achieving state-of-the-art performance for TAD.
@@
Multi-representation fusion design (Fig. 2, c).  Fusion-based models, which currently report state-of-the-art performance in TAD, explicitly combine multiple information sources. TTHF [24] augments the CLIP [36] framework with a high-frequency motion encoder and a cross-modal fusion module to align motion features with textual prompts. PromptTAD [35] extends MOVAD [37] by incorporating bounding box prompts via instance- and relation-level attention, enhancing object-centric anomaly localization. ISCRTAD [25] integrates agent features (e.g., appearance, trajectory, depth) using graph-based modeling and fuses them with scene context through contrastive multimodal alignment for robust anomaly detection.
@@
In this work, we follow the simple encoder-only design and explore whether advanced pre-training can compensate for the lack of task-specific architectural inductive biases. We demonstrate that, when equipped with rich priors from large-scale self-supervised pre-training, such models can achieve state-of-the-art performance, while remaining architecturally simple and highly efficient.
@@ 2.2 Video Foundation Models
2021 ViViT [1] 1 FSL Classification Class labels
2022 VideoMAE [44] 1 SSL Masked autoencoding Video frame pixels
2023 MVD [51] 1 SSL Masked feature distillation High-level features of VideoMAE and ImageMAE teachers
2023 VideoMAE2 [50] 1 SSL Dual masked autoencoding Video frame pixels
2 FSL Classification Class labels
3 FSL Logit distillation Logits of larger VideoMAE2 after stage 2
2023 UMT [22] 1 WSL Unmasked feature distillation Features of a vision-language encoder
2024 InternVideo2 [52] 1 WSL+SSL Unmasked feature distillation Features of VideoMAE2 and a vision-language encoder
2 WSL Contrastive Video-text and audio-text pairs
3 WSL+SSL Feature distillation Features of InternVideo2 after stage 2 across multiple depths
2025 SMILE [43] 1,2 WSL Masked feature distillation Features of a vision-language encoder
@@ 3 WSL+SSL Feature distillation Features of InternVideo2 after stage 2 across multiple depths
While early 3D CNN-based architectures can be referred to as foundation models [31], the term foundation model today typically refers to Transformer-based [46] models that leverage large-scale pre-training. For vision, these models typically adopt the Vision Transformer [9] (ViT) architecture. Although other architectures such as recurrent, hybrid, and state-space models are active research areas [23, 53, 34], ViT-based ViFMs are currently the dominant paradigm due to their strong performance, scalability, versatile pre-training strategies, and widespread optimization support (e.g., FlashAttention [7], optimized libraries, hardware acceleration) designed around the plain Transformer architecture. These qualities, along with the growing availability of pre-trained Video ViTs, make them particularly promising for tasks like TAD, where generalization, robustness, and efficiency are critical.
Unlike convolutional architectures, which impose strong spatial and temporal inductive biases by design, ViTs must learn such priors directly from data during pre-training [9]. As a result, their downstream performance depends heavily on the scale and quality of the pre-training dataset, as well as the choice of training objective [56]. Pre-training strategies differ in supervision strength and scalability:
Fully-supervised learning (FSL) methods rely on manually annotated labels, providing precise semantic guidance but limited scalability. First Video ViTs exploited supervised classification as their pre-training method. ViViT [1] adopted the ViT architecture to video by introducing spatio-temporal 3D cubes called tubelets instead of 2D patches used for images. While this demonstrated that attention-based models can handle video inputs, the method struggled to balance accuracy and efficiency.
Weakly-supervised learning (WSL) methods use natural language or metadata as training signals. Though less curated, they offer rich semantic structure and broader concept coverage from web-scale data. HowTo100M [32] introduced large-scale WSL pre-training by aligning video instructions with narrations using contrastive loss, which MIL-NCE [33] extended with Multiple Instance Learning (MIL) to handle temporally imprecise supervision. UMT [22] and SMILE [43] leverage feature distillation from a vision-language model; UMT adopts an unmasked token alignment, while SMILE introduces motion-guided masking for better semantic and temporal understanding.
Self-supervised learning (SSL) methods learn from the data itself without any annotations, enabling large-scale pre-training and highly transferable representations. Unlike FSL and WSL, which introduce annotation noise and bias, SSL provides denser and more unbiased training signals [4, 14]. VideoMAE [44] adopted masked autoencoding (MAE) [14], a type of Masked Video Modeling (MVM), as an effective and efficient self-supervised pre-training strategy for plain video ViTs. Its tube masking, when applied to a large fraction of input patches, forces the model to infer spatio-temporal structure from limited visible content. This approach yielded strong results while maintaining architectural simplicity and has since inspired a series of ViT-based Video Foundation Models (ViFMs) that employ self-supervised pre-training [51, 50, 41, 15, 38, 43].
@@
To our knowledge, we are the first to research which type of video pre-training is most effective for the TAD task, and hypothesize that self-supervised pre-training with an MVM objective, with its emphasis on learning patch-dense and temporally-aware representations, is particularly well-suited for this task.
@@ 3 Methodology
We fine-tune multiple Video Foundation Models (ViFMs) for TAD. We follow the encoder-only design and attach a single linear layer as a classification head to the output of the encoder. This minimal design ensures that performance primarily reflects the effectiveness of the ViFM backbone in capturing patterns relevant for traffic anomaly detection.
Given a video input $X_{t}\in\mathbb{R}^{T\times H\times W\times 3}$ at time $t$, we extract spatio-temporal feature vector $F_{t}\in\mathbb{R}^{E}$ using the encoder (ViFM), where $T$, $H$, and $W$ are the number of frames, height, and width, and $E$ is the feature dimension. Then, a linear layer maps $F_{t}$ to classification logits $L_{t}\in\mathbb{R}^{C}$, with $C=2$. We fine-tune using cross-entropy loss and apply softmax during inference to obtain anomaly class probability as the anomaly score, which can then be binarized with a decision threshold (typically 0.5). More details in Appendix A.
We investigate (i) whether a simple Video ViT model, pre-trained at scale, can achieve state-of-the-art performance on TAD, (ii) whether better ViFMs for general video recognition are also better for TAD, and what type of pre-training is more effective, and, finally, (iii) whether small-scale domain-adaptive pre-training (DAPT) is feasible and effective for adapting Video ViTs to the driving domain.